OpenStudy (anonymous):

At what value of X does the minimum or maximum occur for the function. F(x)= 2 (x+2)^2+8 A.) Minimum occurs at x=8 B.)Maximum occurs at X=-2 C.)Maximum occurs at x=8 D.) Minimum occurs at x=-2

OpenStudy (anonymous):

For this, we need to calculate the derivative: f(x)=2(x+2)² + 8 f'(x)=4x + 8 To determine where the minimum or maximum is, we solve f'(x)=0 4x+8=0 4x=-8 x=-2 Now, does f(x) have a maximum or a minimum?

OpenStudy (anonymous):

maximum?

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

f(x)=2(x+2)² +8 = 2(x²+4x+4)+8=2x²+8x+16 Now, a=2, so f(x) has a minimum. Look: y=x² has a minimum, and a=1, which is positive.

OpenStudy (anonymous):

So the answer is D

OpenStudy (anonymous):

oh okay so it's minimum... duhhh and that would make it D

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

That's my duty as mathematician.